OpenStudy (anonymous):

SOLVE EQUATION 8/x=x/x-2

OpenStudy (anonymous):

ok so you are solving for x which means you have to get x on one side, do you know how to do this

OpenStudy (anonymous):

@xoxox123

OpenStudy (anonymous):

ok so sorry, i haven't done this for a while, for all items, you would want to do the opposite of it on both sides (example division opposite is multiplication)

OpenStudy (abdullahm):

\(\large\bf \frac{8}{x} =\frac{x}{x-2}\) We need to use this: \(\large \bf \frac{a}{b}=\frac{c}{d}\) becomes \(\large\bf ad = bc\) This is called cross multiplying.

OpenStudy (abdullahm):

There has to be an equal sign 8(x-2) = x(x) Solve for x

OpenStudy (abdullahm):

Oh, my bad. We need to set it equal to 0 and then factor it. :)

OpenStudy (anonymous):

\[x ^{2}-8x+16=0\] \[(x-4)^{2} =0\] Answer is x = 4